OpenStudy (anonymous):

The square of the difference of two numbers equals the sum of their squares less twice their products.

OpenStudy (anonymous):

(x-y)^2 = (xy)^2 - (x^2 + y^2)

OpenStudy (anonymous):

Translate this almost literally word by word: The square of: ()^2 The difference of 2 numbers: x-y Equals: = the sum of their squares: x^2 + y^2 less twice their products: -2xy So (x-y)^2 = x^2 + y^2 -2xy

OpenStudy (anonymous):

Thks

OpenStudy (anonymous):

kumarsajan's answer is incorrect I believe